Grouse Corporation, a calendar year C corporation, had taxable income of $1,400,000, $1,200,000, and $700,000 for 2017, 2018, and 2019, respectively. Grouse has taxable income of $1,600,000 for 2020.

What are Grouse Corporation's minimum required estimated tax payments for 2020?

April 15, 2020: $____

June 15, 2020: $ _____

September 15,2020: $ _____

December 15, 2020: $ _______

Total: $ ______
Answer #1

Estimated tax payment for 2020 = taxable income*21% = 1600000*21% = $336000

Estimated tax payment for each installment = 336000/4 = $84000

Grouse Corporation's minimum required estimated tax payments for 2020:

April 15, 2020: $84000

June 15, 2020: $84000

September 15,2020: $84000

December 15, 2020: $84000

Total: $336000